OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 5t. Find the instantaneous velocity at t = 4 by finding the derivative.

OpenStudy (irishboy123):

\[{ds(t) \over dt} = v(t) = { d\over dt}(-9 - 5t)\] \[{d \over dt} (-9) = ??\] \[{d \over dt} (-5t) = ??\]

OpenStudy (anonymous):

what? sorry that is confusing me

OpenStudy (irishboy123):

ok you are asked to find the derivative of: \(s(t) = -9 - 5t\) that will give you an equation for the velocity \(v(t)\) of the object. how do you propose to do that?

OpenStudy (anonymous):

v (t) = ds/dt?

OpenStudy (wmj259):

That is correct. The velocity is the derivative of the position function.

OpenStudy (wmj259):

If you plug in a specific time into the velocity function you find the INSTANTANEOUS SPEED.

OpenStudy (anonymous):

So if i put v (9) = ds/d(9)

OpenStudy (anonymous):

i mean 4

OpenStudy (irishboy123):

i strongly suggest you do the derivative first. then you will see :p

OpenStudy (anonymous):

Yeah see the problem is I dont know how to do any of this, like for this lesson i missed it in school, and im doing a review right now for the quiz tomorrow and i dont understand this at all

OpenStudy (wmj259):

@IrishBoy123, Yes that would make your night much easier. By taking the derivative of the position function you get the velocity function. But in this case its not much of a velocity FUNCTION then it is a velocity CONSTANT.

OpenStudy (anonymous):

sO IT WOULD BE a velocity constant? whats that

OpenStudy (anonymous):

@Loser66 @satellite73

OpenStudy (wmj259):

Yes, But if you take the derivative of the position function, what do you get?

OpenStudy (anonymous):

idk? how would i find that